OpenStudy (anonymous):

A drawer contains 4 red socks, 3 white socks, and 3 blue socks. Without looking, you select a sock at random, replace it, and select a second sock at random. What is the probability that the first sock is blue and the second sock is red?

mathslover (mathslover):

Ok !!! so first of all .... probability of each event .. case -1 .. probability of pulling blue sock

mathslover (mathslover):

we will call this as s(b) ok ?

OpenStudy (anonymous):

okay well is it 7/20?

mathslover (mathslover):

no ...wait !!! think it again

mathslover (mathslover):

no. of blue socks / total no. of socks .. can u find this ?

OpenStudy (anonymous):

so 3 divided by 10...

mathslover (mathslover):

yes right

OpenStudy (anonymous):

0.3...

mathslover (mathslover):

s(b)=\(\large{\frac{3}{10}}\)

mathslover (mathslover):

now we will be moving on to case-2 that is of finding s(red)=s(r) no. of red socks/total no. of socks

mathslover (mathslover):

can u find s(r) ?

OpenStudy (anonymous):

4/10

mathslover (mathslover):

very correct ! good

mathslover (mathslover):

s(b and r)=s(b)*s*(r) calculate this now

mathslover (mathslover):

put the value of s(b)=3/10 and that of s(r) =4/10 multiply both s(b)*s(r) .. what r u getting ?

mathslover (mathslover):

wait ..think it again kmstevens10 ... what is 3/10 *4/10 ?

mathslover (mathslover):

\[\large{\frac{3}{10}*\frac{4}{10}}\] try to calculate this again

OpenStudy (anonymous):

12/100

mathslover (mathslover):

right !!! so this is ur answer

OpenStudy (anonymous):

that is not one of the answers they give me

mathslover (mathslover):

What r the options

OpenStudy (anonymous):

a. 3/5 b. 7/20 c. 3/25 d. 7/100

mathslover (mathslover):

u can simplify 12/100 as 3/25

OpenStudy (anonymous):

thank you.

mathslover (mathslover):

\[\large{\frac{12}{100}=\frac{4*3}{4*25}=\frac{\cancel4 *3}{\cancel4*25}=\frac{3}{25}}\]

OpenStudy (anonymous):

Thank you so much I get it now the next one can you help me with step by step that was a huge help ....

OpenStudy (anonymous):

Two urns each contain green balls and blue balls. Urn I contains 4 green balls and 6 blue balls, and Urn II contains 6 green balls and 2 blue balls. A ball is drawn at random from each urn. What is the probability that both balls are blue?

mathslover (mathslover):

again there are 2 cases : urn 1 and urn 2 urn 1 : total no. of balls = 10 no. of blue balls = 5

mathslover (mathslover):

can u find s(urn1) ? for blue balls \[\large{\frac{\textbf{no. of blue balls}}{\textbf{total no. of balls}}}\] in urn 1

OpenStudy (anonymous):

18 balls correct?

mathslover (mathslover):

no ! in urn 1 .. no. of blue balls = 5 . total no. of balls = 10 hence p(b)=5/10

OpenStudy (anonymous):

ok so 5/10

mathslover (mathslover):

yes now in urn II : 2/10 =s(b) p(b)*s(b)=5/10 * 2/10

OpenStudy (anonymous):

the whole problem is simplified down to 1/10

mathslover (mathslover):

yes !!

OpenStudy (anonymous):

thank you again!